mikewwe13:

Graph f(x) = 2^x + 1 and g(x) = −x + 7 on the same coordinate plane. What is the solution to the equation f(x) = g(x) ? Enter your answer in the box. x = _______

mikewwe13:

@Vocaloid

Vocaloid:

have you tried graphing these yet?

mikewwe13:

not yet but now

mikewwe13:

1 attachment
Vocaloid:

well done so what is the x-value of the intersection?

mikewwe13:

7

Vocaloid:



Vocaloid:



mikewwe13:

2

Vocaloid:

the solution is x = 2

mikewwe13:

SORRY I didn't see it

mikewwe13:

What is the solution to f(x)=g(x) ? Select each correct answer. x = −2 x = −1 x = 0

1 attachment
Vocaloid:

hint: which two rows of the table f(x) do g(x) have the same value?

mikewwe13:

-1

Vocaloid:

good but there's one more x value where f(x) = g(x)

mikewwe13:

and 1

Vocaloid:

1 is not one of the answer choices please look carefully

mikewwe13:

and 0

mikewwe13:

my computer wasn't acting right

Vocaloid:

good, -1 and 0 = your answers

mikewwe13:

Let f(x) = 2x and g(x) = 2^x. Graph the functions on the same coordinate plane. What are the solutions to the equation f(x) = g(x) ? Enter your answers in the boxes. x = ______ or x = _________

Vocaloid:

have you tried graphing this?

mikewwe13:

1 attachment
mikewwe13:

here it is